OpenStudy (anonymous):

what will be the period of the function f(t)=(pi/6)*t + 6sin (pi/3)t ??

OpenStudy (tanya123):

f(t) = (pi t)/6+3 sqrt(3) t

OpenStudy (anonymous):

attached the graph of the function...

OpenStudy (anonymous):

the period is \[T=\frac{2\pi}{\frac{\pi}{3}}=2\pi \times \frac{3}{\pi}=6\]

OpenStudy (anonymous):

function is not periodic :D

OpenStudy (anonymous):

@orion 1213 is the function periodic in t .... i.e along the t axis ????

OpenStudy (anonymous):

no it's not at t-axis but periodic at line \[\frac{\pi}{6}\times t\], the 1st term of the function...